If the license plates in a Karachi consist of three letters followed by three numbers, what is the probability that a randomly generated plate ends in the numbers “786”? Consider that repetitions are allowed and order matters.
1
Expert's answer
2020-06-15T11:33:24-0400

we have 10 options for each position - from 0 to 9

we need 7 on first number-position, P = 1/10=0.1; 8 on second number-position, P = 1/10=0.1; 6 on third number-position, P = 1/10=0.1

So, P=0.1*0.1*0.1=0.001